OpenStudy (anonymous):

Without a calculator determine the exact value of tan 7pi/4

OpenStudy (anonymous):

do you have a unit circle?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

\[\tan \frac{ 7\pi }{4 }=\tan \left( 2\pi-\frac{ \pi }{4 } \right)=-\tan \frac{ \pi }{4 }= -1\]

OpenStudy (anonymous):